# FaceEditTalker: Interactive Talking Head Generation with Facial Attribute Editing

### v1

@@ 4.3 Analysis and Ablation Study
Table 4: Quantitative Metrics for Ablation Study of Facial Landmark Feature Extractor
@@
Original Video 7.359 7.586 0.000 0 0
Multi-layer Convolution 12.534 6.562 7.000 10.468 7.892
Multi-scale Strategy 9.300 5.344 3.000 4.532 5.246
Multi-scale & Cross-Attention 8.145 6.288 0.000 3.354 3.465
[…]
Effectiveness of the Facial Landmark Feature Extractor. We conducted ablation experiments on two datasets to validate the multi-scale strategy and cross-attention mechanism in the facial landmark feature extractor. The multi-scale strategy improved fine-grained movements like lip and eye motions, while the cross-attention mechanism enhanced all metrics, particularly AVOffset and M-LMD, achieving better consistency with the original video. Furthermore, testing the Edited Semantic Encoding showed minimal impact on evaluation metrics.
@@
Linear Distribution of Attributes in Latent Space. To verify the linear distribution of target attributes in latent space, we first visualized the latent space using Principal Component Analysis (PCA). The results show that in the PCA space, samples of different attributes exhibit clear separation along the principal component directions, proving that semantic latent variables exhibit a linear distribution in the latent space. This validates the rationality of linear operations in the semantic space and supports the classifier’s ability to distinguish between different attribute categories. Additionally, to demonstrate the interpolability of high-level semantic features across different attributes, we generated videos by interpolating features from two different speakers’ images. The interpolation results appeared very natural, showing good separation of attributes.
5 Conclusion and Limitations
@@
Conclusion. In summary, we have introduced a novel framework for editable talking face generation that significantly enhances both the realism and controllability of facial animations. By leveraging a combination of disentangled latent representations and fine-grained audio-visual alignment, our method enables intuitive editing capabilities such as face editing and lip-sync correction. Extensive experiments on multiple benchmarks demonstrate that our approach not only achieves superior performance compared to existing state-of-the-art methods but also allows for diverse and personalized talking face generation. We believe this framework opens up promising directions for future research in personalized avatars, virtual assistants, and digital human synthesis.
@@
Limitations. Despite the success of our framework, we also recognize some limitations during the exploration. First, while demonstrating good generalization ability, performance might degrade in environments with highly diverse identities, complex head movements, or challenging conditions that are not well-represented in the training data. Second, the diffusion-based generation process, although yielding high-quality results, currently incurs significant computational costs, limiting its application in real-time scenarios. Third, currently our method can only support the editing of 40 preset attributes. In the future, we hope to leverage the capabilities of models like CLIP to achieve flexible face attribute editing through arbitrary text descriptions.
Ethical Considerations. Our talking head generation method is strictly limited to academic research and will not be applied to unethical domains like fraud or misinformation. The generated model will be shared with the Deepfake detection community to support identification research, ensuring responsible development of this technology.
6 Acknowledgments
@@
The work was jointly supported by the National Science and Technology Major Project under grant No. 2022ZD0117103, the National Natural Science Foundations of China under grant No. 62272364, 62472342, the Provincial Key Research and Development Program of Shaanxi under grant No. 2024GH-ZDXM-47, the Research Project on Higher Education Teaching Reform of Shaanxi Province under Grant No. 23JG003.
Appendix A Pseudo-code for this method
A.1 Training and Inference Process Description
The training phase consists of three main modules:
1. Joint Training of Semantic Encoder and Random Encoder: Face images are encoded to extract high-level semantic and random features, optimized using a diffusion model.
2. Training the Image Semantic Linear Classifier: Attribute variation directions are learned in the high-level semantic space, enabling precise attribute classification.
3. Training the Latent Space Diffusion Model: Multi-scale facial motion features, semantic encodings of reference images, and audio-driven information are integrated to generate high-quality, identity-consistent talking face videos.
In the inference phase, the model uses the input audio sequence, reference image, and attribute information to generate facial motion features, which are processed by the diffusion model to produce high-quality, editable talking face videos.
A.2 Training Stage
Algorithm 1 Joint Training of Semantic Encoder and Random Encoder
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
[…]
5:         $z_{\text{sem}}=\text{SemanticEncode}(x_{i})$triangleright Semantic Encoder Forward Pass
6:         $t=\text{RandomTimeStep}()$
7:         $z_{0}=\text{Encode}(x_{i})$
8:         $x_{t}=\text{AddNoise}(z_{0})$
9:         $\hat{x}=\hat{q}(x_{t},z_{\text{sem}})$triangleright Model Forward Pass
10:         $e=\text{ComputeTarget}(z_{t})$triangleright Compute Target
11:         $L=\text{MSE}(e,\hat{x})$triangleright Loss Calculation
12:         Backpropagate and update parameters
13:     end for
14:     Save model at the end of each epoch
15:end for
Algorithm 2 Training the Image Semantic Linear Classifier
1:Input: Face image set, attribute labels, learning rate
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
4:     for each attribute $a_{i}\in\text{Attr}(a_{1},a_{2},\dots,a_{n})$ do
5:         $z_{\text{sem}}=\text{SemanticEncode}(x_{i})$
6:         $y_{\text{labels}}=\text{GetAttributeLabels}(a_{i},\text{FFHQ}(image))$
[…]
8:         $b=\text{InitializeBias}()$
[…]
10:              for each $z_{\text{sem}}$ do
11:                  $y_{\hat{t}}=\text{Sigmoid}(W^{T}z_{\text{sem}}+b)$triangleright Forward Pass
12:                  $L=\text{CrossEntropyLoss}(y_{\text{labels}},y_{\hat{t}})$triangleright Cross-Entropy Loss
13:                  Backpropagate and update parameters
14:              end for
15:         end for
16:     end for
17:     Save weight vectors and attribute-label pairs
18:end for
Algorithm 3 Training the Latent Space Diffusion Model Generation Network
1:Input: Face image set, reference image set, learning rate
2:Output: Diffusion model parameters
3:for each epoch do
4:     for each batch $x_{r},x_{\text{ref}},l_{r},l_{\text{ref}}$ do
5:         $z_{0}=\text{Encode}(x_{r})$
6:         $z_{t}=\text{AddNoise}(z_{0})$
7:         $t=\text{RandomTimeStep}()$
8:         $z_{\text{ref}}=\text{SemanticEncode}(x_{\text{ref}})$
9:         $z_{1}=\text{PoseGuided}(l_{r},l_{\text{ref}})$
10:         $y_{\hat{t}}=\hat{q}(z_{1},z_{t},z_{\text{ref}})$
11:         $e=\text{ComputeTarget}(z_{t})$
12:         $L=\text{MSE}(e,y_{\hat{t}})$triangleright MSE Loss
13:         Backpropagate and update parameters
14:     end for
15:     Save model at the end of each epoch
16:end for
A.3 Inference Stage
Algorithm 4 Based on Linear Space Facial High-Level Semantic Feature Editing Module (Inference Process)
1:Input: Reference image $x_{ref}$, audio sequence $\mathcal{S}_{audio}$, attribute labels $(y,w)$, attribute editing magnitude, diffusion time steps steps
2:Output: Speaker video frame sequence $S_{video}$
3:Data Preprocessing
4:Use Wav2Vec to extract audio feature sequence from audio sequence $\mathcal{S}_{audio}$
5:$F^{1:T}=F(f_{1},f_{2},...,f_{t})=Wav2Vec(S_{audio})$
6:$x_{mesh}=MediaPipeMesh3D(x_{ref})$
7:$M^{1:T}=Audio2Mesh(F^{1:T},x_{mesh})$
8:$P^{1:T}=Audio2Pose(F^{1:T})$
9:$L^{1:T}=PerspectiveProjection(M^{1:T},P^{1:T})$
10:$l_{ref}=MediaPipeFaceLandmarker(x_{ref})$
11:$z_{l}^{1:T}=PoseGuider(L^{1:T},l_{ref})$
12:$z_{sem}=SemanticEncode(x_{ref})$
13:if a is not None then
14:     $z_{sem}=z_{sem}+\alpha w$
15:end if
16:$S_{video}=DiffusionModel(z_{sem},steps,z_{l}^{1:T})$
Appendix B Experimental parameter settings
Our methods was trained on two A100 GPUs. The first and second stages were trained for 100 hours, and the third stage was trained for 160 hours. The main parameters are shown in Table 5.
Table 5: Experimental Parameter Settings
Parameters Value/Range
Random Seed 0
Image Size 512*512
Batch Size 16
Learning Rate 0.0001
Training Epochs 20000
Embedding Layer Channels 512
Diffusion Timesteps 1000
Appendix C More Qualitative Evaluation
More qualitative evaluation can be seen in Figures 7-14, which include 40 attribute editing effects of two celebrities, and the dynamic video can be found in the attachment.
Figure 7: Accessories & Makeup of Jay Chou
Figure 8: Accessories & Makeup of Michelle Yeoh
Figure 9: Facial State of Jay Chou
Figure 10: Facial State of Michelle Yeoh
Figure 11: Facial Structure of Jay Chou
Figure 12: Facial Structure of Michelle Yeoh
Figure 13: Hair Features of Jay Chou
Figure 14: Hair Features of Michelle Yeoh

### v2

@@ V-C Analysis and Ablation Study
TABLE V: Quantitative Metrics for Ablation Study of Facial Landmark Feature Extractor.
@@
I 7.359 7.586 0.000 0.000 0.000
II 12.534 6.562 7.000 10.468 7.892
III 9.300 5.344 3.000 4.532 5.246
IV 8.145 6.288 0.000 3.354 3.465
V 8.484 6.894 0.000 3.301 3.566
Linear Distribution of Attributes in Latent Space. To verify the linear distribution of target attributes in latent space, we first visualized the latent space using Principal Component Analysis (PCA). Fig.6 shows that in the PCA space, samples of different attributes exhibit clear and meaningful separation along the principal component directions, proving that semantic latent variables exhibit a linear distribution in the latent space. This validates the rationality of linear operations in the semantic space and supports the classifier’s ability to distinguish between different attribute categories. Additionally, to demonstrate the interpolation capability of high-level semantic features across different attributes, we generated videos by interpolating features from two different speakers’ images. The interpolation results appeared very natural in Fig.7 , showing good separation of attributes.
@@
VI Conclusion
VI-A Conclusion
@@
In summary, we have introduced a novel framework for editable talking face generation that significantly enhances both the realism and controllability of facial animations. By leveraging a combination of disentangled latent representations and fine-grained audio-visual alignment, our method enables intuitive editing capabilities such as face editing and lip-sync correction. Extensive experiments on representative baseline methods demonstrate that our approach not only achieves superior performance compared to existing representative methods, but also allows for diverse and personalized talking face generation. We believe this framework opens up promising directions for future research in personalized avatars, virtual assistants, and digital human synthesis.
VI-B Limitation
@@
Despite the success of our framework, we acknowledge several limitations. First, although it demonstrates strong generalization ability, performance may degrade in scenarios involving highly diverse identities, complex head movements, or challenging conditions insufficiently represented in the training data. Second, the diffusion-based generation process, while producing high-quality results, incurs substantial computational costs, limiting its applicability in real-time settings. Third, our method currently supports only 40 predefined facial attributes and does not allow personalized image generation conditioned on arbitrary text prompts. Moreover, while directly extracting facial keypoints from audio improves lip-sync accuracy, it often leads to rigid or overly uniform facial expressions, reducing the realism of generated videos—as reflected in the user study results (Table IV). In the future, we aim to leverage models such as CLIP to enable flexible facial attribute editing through arbitrary text descriptions and to explore strategies that jointly preserve accurate lip synchronization and natural facial dynamics.
VI-C Ethical Consideration
We carefully consider the ethical implications of our work. Methods such as FaceEditTalker, which enable facial attribute editing, entail inherent risks, including unauthorized use of personal likeness and the spread of misleading content. To address these concerns, we restrict our framework strictly to academic research and prohibit any fraudulent or deceptive applications. The trained model and supporting detection resources will be shared only with the deepfake detection research community to strengthen identification efforts and promote the responsible advancement of this technology.
VII Acknowledgments
@@
The work was jointly supported by the National Natural Science Foundations of China under grants No. 62272364, 62472342, the Provincial Key Research and Development Program of Shaanxi under grant No. 2024GHZDXM-47, the Research Project on Higher Education Teaching Reform of Shaanxi Province under Grant No. 23JG003.